OpenStudy (anonymous):

-2/3t-3=-5 a. t= -1/3 b. t= 1/3 c. t= 2/3 d. t= 3

OpenStudy (anonymous):

well madam, :  :)

OpenStudy (senkai):

-2/3t - 3 = -5 Add 3 to both sides. So the equation now would be: -2/3t = -2 For -2/3, you need to flip the fraction to 3/-2, so you can cancel them out and get t by itself. Multiply 3/-2 to both sides. t= -2(3/-2) D. t=3.

OpenStudy (anonymous):